Suppose the IS curve is Y = 2500-125i and the LM is Y = 1000+25i. Use Cramer’s rule to determine i* and Y*. Now suppose increased spending by $100m. Determine the extent of crowding out of private investment and the increase in Ms required to dampen any crowding-out effect. You are given the following parameters: αG = 2.5, h = 65, k = 0.5, b = 50.
1
Expert's answer
2020-05-11T19:24:34-0400

According to Cramer’s rule:

IS = LM,

2500 - 125i = 1000 + 25i,

150i = 1500,

i* = 10,

Y* = 1000 + 25×10 = 1250.

The fiscal policy such as increase in spending by $100m will be crowded out, so the increase in private investment will be lower than 100m and, that's why such monetary policy as the increase in Ms is required to dampen any crowding-out effect.